OpenStudy (anonymous):

General solution of (1+x^2)dy/dx + 2xy = 3x^2?

OpenStudy (anonymous):

What are you solving for x, dy/dx?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

I don't get it. Is this an algebra exercise? Solve for y?

OpenStudy (anonymous):

y in terms of x (you need some kind of integration to get rid of the dy/dx).

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Everybody seems very hot on integration methods so I thought we would step up a little:-)

OpenStudy (anonymous):

divide by (1 + x^2) on both sides

OpenStudy (anonymous):

this is homogenous equation and should be solved by putting y = vx

OpenStudy (anonymous):

dy/dx + 2xy/(1+x^2) = 3x^2 / (1+x^2)

OpenStudy (anonymous):

Integrating Factor = \[e^{\int\limits_{}^{}[2xdx/(1+x^2)]}\]

OpenStudy (anonymous):

multiply by IF on both sides

OpenStudy (anonymous):

(1 + x^2)y = 3x^2dx

OpenStudy (anonymous):

Making progress...2x is derivative of x^2+1 which should remind you of the Product Rule.

OpenStudy (anonymous):

ur testing me? i already know how to do this

OpenStudy (anonymous):

y = x^3/ (1+x^2) + c

OpenStudy (anonymous):

If you know, let others try...

OpenStudy (anonymous):

u should put up a signpost saying "TRAINING EXERCISE" then, coz i thot u needed the explanation

OpenStudy (anonymous):

Maybe the others would like to see the explanation?

OpenStudy (anonymous):

i did it on top...its a linear differential eqn..

OpenStudy (anonymous):

once you divide by 1+x^2 on both sides

OpenStudy (anonymous):

Why does it work?

OpenStudy (anonymous):

why does what work?

OpenStudy (anonymous):

he divided it by 1 + x^2 to get it in linear form, which is: dy/dx + P(x)*y = Q(x)

OpenStudy (anonymous):

now u got it?

OpenStudy (anonymous):

whers he?

OpenStudy (anonymous):

Sorry, system went down...

OpenStudy (anonymous):

Yes, I have contrived this example in order to show how it works...

OpenStudy (anonymous):

I do apologize for any confusion...

OpenStudy (anonymous):

D(1+x^2)y = (1+x^2)dy/dx + 2xy (product Rule) and the RHS is same as LHS of original eq. So D(1+x^2)y) = 3x^2 Now lhs is just the derivative of (1+x^2)y so you can apply direct integration.. OK, contrived, so how to find the IF when it is not so convenient?

OpenStudy (anonymous):

I can type out the diffeq solution if you want. Using the method of integrating factors.

OpenStudy (anonymous):

Well thanks for that, following the method is OK of course, I guess I am just one of those that likes to see the "why" of the method and I was wondering if others thought the same. Perhaps unfashionable? Or maybe this is just not the right sort of forum for that sort of thing as I haven't seen anyone put up a "TRAINING EXERCISE" notice to their questions.

OpenStudy (anonymous):

I can get out my diffeq book and derive for you the method of integrating factors :P But you can look that up of course :)

OpenStudy (anonymous):

We seem to be talking slightly at cross purposes, I already have the derivation, the question was designed to lead up to it (but we never got that far.-)

OpenStudy (anonymous):

are you a teacher, estudier?

OpenStudy (anonymous):

I was, once upon a time, I am retired now.

OpenStudy (anonymous):

much respect!

OpenStudy (anonymous):

Are there other forums maybe like this but with the possibility for more of a discussion?

OpenStudy (anonymous):

u should ve specified sir, here we think ur a student looking fr answers

OpenStudy (anonymous):

sorry for being rude earlier

OpenStudy (anonymous):

No offence taken, I am learning as well:-)

OpenStudy (anonymous):

r u lookin for the derivation of the integrating factor

OpenStudy (anonymous):

I had in mind to explain it to some of the students but I didn't really know if there were any that wanted to know it, so I thought to come at it indirectly via a series of integration questions.

OpenStudy (anonymous):

The level here seems to be something like A-level/first year Uni and with many below that, is that right?

OpenStudy (anonymous):

yes - thats about right - theres a wide range of levels. By the way i tried solving that diff. equation by putting y = vx and separate the variables but the algebra was heavy - i gave up - might be my age i suppose!

OpenStudy (anonymous):

estudier, thank you for teaching me that the integrating factgor method is derived from the product rule

OpenStudy (anonymous):

jimmy, the equation is not separable, even with y = vx

OpenStudy (anonymous):

yep - u must be right - I mean -even I failed!! lol!

OpenStudy (anonymous):

@jamesm That is the key point, sep var relies on integration by substitution which is equivalent to the chain rule. So it is natural to ask whether you can solve fode deriving from integration by parts (ie the product rule for derivatives).

OpenStudy (anonymous):

I will write something up and leave it here for anyone to look at if they want to.

OpenStudy (anonymous):

great idea

OpenStudy (anonymous):

is the equation is homogenous estuder? seems so to me it is. I guess that not all of them can be solved with y=vx

OpenStudy (anonymous):

Linear fode if:- dy/dx + g(x) y = h(x) Linear fode homogenous if h(x) = 0 for all x (else inhomogenous).

OpenStudy (anonymous):

Let me do writeup, then questions (if any).

OpenStudy (anonymous):

To recap, we got the answer for the original question via seeing that lhs was like p dy/dx + dp/dx y (p = 1 + x^2) and reexpressing it with the Product rule as d p(y)/dx. If we rewrite the original as dy/dx + 2xy/(1+x^2) = 3x^2/(1+x^2), you can see it is the multiplying through (which was already done in the original question) by (1+x^2) that allowed the quick result because it allowed direct integration. So how to find this factor usually? Well, it has to be such that when you multiply on the lhs it looks like p dy/dx + p(2x/(1+x^2))y and also like p dy/dx + dp/dx y So if you compare these two then you can see that p is a particular solution of dp/dx = (2x/(1+x^2)) p which is homogeneous linear fode solvable by sep var to give |p| = exp (int(2x/(1+x^2)) dx). This can be generalized for any linear fode of form dy/dx + g(x)y = h(x) whereby g(x) replaces the 2x/(1-x^2) of our original question. And that is where the method comes from, exp your g(x), multiply through by p(x), rearrange and direct integrate (although just as with sep var, it might not be possible to perform the final integration).

OpenStudy (anonymous):

If people get it, we can go on to to sode's...

OpenStudy (anonymous):

yes - I follow you ok . I didn't rewalise how rusty I was with fodes!